how much 20% antifreeze solution must be mixed with a 40% antifreeze solution to produce 25 gallons of a 32% disinfectant solution? round to the nearest gallon.
? gallons 20% solution, ? gallons 40% solution

Answer by checkley77(12569)

.40x+.20(25-x)=.32*25
.40x+5-.20x=8
.20x=8-5
.20x=3
x=3/-20
x=15 gal. of 40% is used.
25-15=10 gal. of 20% is used.
Proof:
.40*15+.20*10=8
6+2=8
8=8
10 gallons 20% solution, 15 gallons 40% solution